OpenStudy (anonymous):

Connor and Matt walk a 12 mile couse as part of a fitness program. Matt walks 1 mi/hr faster than Connor, and it takes him 1 hour less than Connor to complete the course. How long does it take Connor to complete the course? (work too please) (fyi answer is 5 hours)

OpenStudy (anonymous):

Let c equal Connor's rate in miles per hour. c*t = 12 For Matt: (c + 1) (t - 1) = 12 The solution to these simultaneous equations is: c = 3, t = 4 For Connor: 3*4=12, therefore, Connor takes four hours to complete the course at 3 mph. (3+1)(4-1)=12, therefore, Matt takes three hours to complete the course at 4 mph, 1 mph faster than and 1 hour less than Connor. I could be wrong but "(fyi answer is 5 hours)" is not supported by the solutions.